Hi please help me. Correct me if I'm wrong plz. The question is
(6t^-7)/ (2t^9)
is the answer 3t^-16 or 3t^2
: Hi please help me. Correct me if I'm wrong plz. The question is
(6t^-7)/ (2t^9)
is the answer 3t^-16 or 3t^2

Answer by KnightOwlTutor(291)

(6t^-7)/ (2t^9)

(6t^-7)=(6/t^7)*(1/2t^9)=3t^-16
You can divde 6 by 2 and get 3 and add the negative exponents together to get -16
3/t^7+t^9=3/t^16=3t^-16